# RIZE: Regularized Imitation Learning via Distributional Reinforcement Learning

### v2

@@ B.2 Implementation Details
Environment $\alpha$ (3 / 10) $c$ lr $\pi$ lr $Z$ $\lambda^{\pi_{E}}$ lr $\lambda^{\pi_{E}}$ $\lambda^{\pi}$ lr $\lambda^{\pi}$ (3 / 10)
Ant-v2 0.05 / 0.10 0.1 $5\times 10^{-5}$ $3\times 10^{-4}$ 10 $1\times 10^{-4}$ 5 $\times 10^{-5} / 1×1$$\times 10^{-4}$
HalfCheetah-v2 0.05 / 0.10 0.1 $5\times 10^{-5}$ $3\times 10^{-4}$ 10 $1\times 10^{-4}$ 5 $\times 10^{-5} / 1×1$$\times 10^{-4}$
Walker2d-v2 0.05 / 0.10 0.1 $5\times 10^{-5}$ $3\times 10^{-4}$ 10 $1\times 10^{-4}$ 5 $\times 10^{-5} / 1×1$$\times 10^{-4}$
Humanoid-v2 0.05 / 0.10 0.5 $1\times 10^{-5}$ $3\times 10^{-4}$ 10 $1\times 10^{-4}$ 5 $5\times 10^{-5}$ / $1\times 10^{-5}$
B.3 Hyperparameter Tuning
We present our analysis and comparison of important hyperparameters utilized in our algorithm. As before, all experiments use five seeds, and we show mean and half a standard deviation of all the seeds. As a challenging task among MuJoCo environments, we only experiment with Ant-v2.
Target Reward $\lambda$
@@ B.3 Hyperparameter Tuning
Selecting appropriate initial values and learning rates for the automatic fine-tuning of $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$ is critical in our approach. First, we observe that a suitable learning rate is essential for the stable training of our imitation learning agent, as illustrated in Figure 7(a). Our findings indicate that $\lambda^{\pi}$ must be optimized very slowly; using larger learning rates can destabilize training and hinder progress. In contrast, $\lambda^{\pi_{E}}$ demonstrates greater resilience when optimized with higher learning rates. Additionally, $\lambda^{\pi_{E}}$ remains robust even with varying initial values. However, as shown in Figure 7(b), failing to select an appropriate initial value for $\lambda^{\pi}$ can negatively impact learning. Overall, Figures 7(a) and 7(b) highlight the need for careful selection of both the learning rate and initial value when optimizing $\lambda^{\pi}$, while $\lambda^{\pi_{E}}$ exhibits considerable robustness in this regard.
@@
Figure 7: Fine-tuning analysis of target parameters. (a) Learning rates: Turquoise represents our method’s primary result with learning rates of $1\mathrm{e}{-4}$ for $\lambda^{\pi_{E}}$ and $1\mathrm{e}{-5}$ for $\lambda^{\pi}$. Orange and blue lines indicate higher learning rates (e.g., $1\mathrm{e}{-3}$) for $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$, respectively. (b) Starting values: Turquoise shows the main result with initial values of 10 for $\lambda^{\pi_{E}}$ and 5 for $\lambda^{\pi}$, while other lines explore different starting values. Three trajectories are used throughout the analysis.
Regularization Coefficient $c$
@@
Entropy Coefficient $\alpha$
We observe that the entropy coefficient is a crucial hyperparameter in inverse reinforcement learning (IRL) problems. As shown in Figure 8(b), IRL methods typically require small values for $\alpha$, a point previously noted (Garg et al., 2021). With expert demonstrations available, an imitation learning (IL) policy does not need to explore for optimal actions, as these are provided by the demonstrations. Consequently, higher values of $\alpha$ can lead to training instability, ultimately resulting in policy collapse.

### v3

@@ Appendix B Implementation Details
Environment $\alpha$ (3 / 10) $c$ lr $\pi$ lr $\lambda^{\pi_{E}}$ lr $\lambda^{\pi}$ (3 / 10)
Ant-v2 0.05 / 0.10 0.1 $5\times 10^{-5}$ $\times 10^{-4} 1×1$$\times 10^{-5}$ / $1\times 10^{-4}$
HalfCheetah-v2 0.05 / 0.10 0.1 $5\times 10^{-5}$ $\times 10^{-4} 1×1$$\times 10^{-5}$ / $1\times 10^{-4}$
Walker2d-v2 0.05 / 0.10 0.1 $5\times 10^{-5}$ $\times 10^{-4} 1×1$$\times 10^{-5}$ / $1\times 10^{-4}$
Hopper-v2 0.20 / 0.20 0.1 $5\times 10^{-5}$ $\times 10^{-4} 1×1$$\times 10^{-4}$ / $1\times 10^{-4}$
Humanoid-v2 0.05 / 0.10 0.5 $\times 10^{-5} 1×1$$\times 10^{-4}$ $5\times 10^{-5}$ / $1\times 10^{-5}$
AdroitHandHammer-v1 0.30 / 0.30 0.1 $3\times 10^{-5}$ $1\times 10^{-4}$ $5\times 10^{-5}$ / $5\times 10^{-5}$
Appendix C Additional Experiments
This section augments the main experimental results with additional experiments, including targeted ablations (critic architecture, regularization design, and loss choice) and a hyperparameter sensitivity analysis.
C.1 Main Results (Extended)
Figure 6 complements the main-text results by presenting learning curves for the three-demonstration regime. Learning curves mirror the 10-demo setting (cf. Figure 3): RIZE and CSIL performs better than the baselines with more stable learning across tasks; CSIL shows strong early progress due to behavior-cloning initialization; IQ-Learn is competitive yet below RIZE; and LSIQ/SQIL trail behind. Task-wise, Humanoid-v2 remains solved only by RIZE, while on Hammer-v1 RIZE and CSIL lead. As expected with fewer demonstrations, normalized returns are lower and variance is higher, but the relative ordering of methods and the sample-efficiency patterns are consistent with the RLiable aggregates reported in the main body (Figure 1).
Figure 6: Learning curves on MuJoCo and Adroit tasks with 3 expert demonstrations. Lines show the mean normalized return across five seeds; shaded regions denote 95% confidence intervals.
C.2 Recovered Reward (Extended)
We complement the main-text analysis with additional plots in the 3–demonstration regime and cross-method comparisons. First, Figure 7 mirrors the main-body analysis for the 10–demo setting: it overlays recovered reward trajectories (expert and policy) with the theoretical bounds implied by our regularizer $\Gamma(R_{Q},\lambda)$ and adaptive targets (Corollary 4.2). As in the main text, the curves remain within the predicted band, and the band’s position/width adapts by task, reflecting how the adaptive targets track and limit expert/policy rewards in practice.
Across all three figures—Figures 9, 8, and 10(a)RIZE and LSIQ keep implicit rewards bounded, whereas SQIL and IQ-Learn exhibit drifting or unbounded rewards on the challenging Humanoid-v2 and Hammer-v1 tasks. In LSIQ, boundedness is likely due to its target clipping; notably, expert-sample rewards concentrate near zero. By contrast, SQIL fails to control reward scales on nearly all tasks, and IQ-Learn is particularly unstable on Humanoid-v2 and Hammer-v1. These observations support the view that adaptive, task-sensitive regularization is effective in maintaining calibrated implicit rewards.
Figure 7: Implicit reward curves for expert and policy samples on MuJoCo and Adroit tasks with 3 expert demonstrations. Each subplot reports the mean across five seeds, with shaded regions showing the 95% confidence interval. Theoretical upper and lower bounds derived in this work are overlaid as separate curves in each subplot.
Figure 8: Implicit reward curves on MuJoCo tasks comparing RIZE with SQIL, IQ-Learn, and LSIQ using 10 expert demonstrations. Lines show the mean across five seeds; shaded regions denote 95% confidence intervals.
Figure 9: Implicit reward curves on MuJoCo tasks comparing RIZE with SQIL, IQ-Learn, and LSIQ using three expert demonstrations. Lines show the mean across five seeds; shaded regions denote 95% confidence intervals.
Figure 10: Adroit Hammer-v1 results: (a) implicit reward curves and (b) value–estimation curves, each reported for 3 and 10 expert demonstrations (n denotes number of trajectories). Lines show the mean across five seeds; shaded regions denote 95% confidence intervals.
C.3 Critic Value-estimation
We examine how critic estimates evolve in RIZE versus baselines. Unlike methods that train point-estimate $Q$-networks, RIZE employs an IQN critic that models the full return distribution $Z(s,a)$ and optimizes losses using its expectation $\mathbb{E}[Z(s,a)]$ (Dabney2018). Because all policies are updated by maximizing state–action values, unbounded estimates can destabilize training and undermine robustness.
Figures 10(b), 11, and 12 show that RIZE and LSIQ maintain bounded values on the challenging Humanoid-v2 and Hammer-v1 tasks, whereas SQIL and IQ-Learn exhibit large, drifting estimates. For LSIQ, boundedness primarily arises from target clipping to [200,200][-200,200], which also explains why expert-sample values cluster near +200+200. Importantly, bounded critics alone do not guarantee expert-level control within our training budgets (3e5 steps for MuJoCo; 5e5 for Adroit): LSIQ remains below RIZE in Figures 3 and 6, indicating that value stability achieved by adaptive reward regularization can translate into better performance than rigid critic clipping with fixed reward targets.
Figure 11: Value-estimation curves on MuJoCo tasks comparing RIZE (IQN critic (Dabney2018)) with SQIL, IQ-Learn, and LSIQ (point-estimate $Q$) using 10 expert demonstrations. Lines are means over five seeds; shaded regions denote 95% confidence intervals.
Figure 12: Value-estimation curves on MuJoCo tasks comparing RIZE (IQN critic (Dabney2018)) with SQIL, IQ-Learn, and LSIQ (point-estimate $Q$) using three expert demonstrations. Lines are means over five seeds; shaded regions denote 95% confidence intervals.
C.4 Ablation on Critic Architecture (Extended)
Figure 13 complements the main-body ablation by examining recovered rewards with theoretical bounds when RIZE uses a point-estimate $Q(s,a)$ critic in place of the IQN return-distribution critic $Z(s,a)$ (cf. Figure 5 for returns, and Figure 7 for bounded rewards under IQN). With the $Q$ critic, implicit rewards are less well constrained in Humanoid-v2 and Hammer-v1 environments: we observe more excursions beyond the predicted band and larger oscillations, in contrast to the IQN setting where rewards remain tightly within the adaptive interval. This comparison suggests that both components contribute to reward stability: the squared-TD regularizer with adaptive targets provides principled bounds, and modeling the return distribution with IQN supplies steadier targets/updates that help keep rewards within those bounds. Together, these effects yield more reliable policy learning and align with the performance gap observed in the return curves.
Figure 13: Implicit reward curves for expert and policy samples on MuJoCo and Adroit tasks using RIZE with a classic $Q(s,a)$ critic. Each subplot shows the mean over five seeds with shaded 95% confidence intervals, using three expert demonstrations. Theoretical upper and lower bounds derived in this work are overlaid as separate curves in each subplot.
C.5 Ablation on Loss Choice (HL-Gauss in LSIQ)
Recent work argues that some of the gains attributed to distributional RL may stem from the loss rather than from modeling the return distribution itself (farebrother2024stop; pmlr-v235-ayoub24a). Following this “loss swapping” view, we replace the mean-squared error (MSE) used in LSIQ’s critic with the HL-Gauss classification loss (farebrother2024stop), which discretizes the value range into bins and trains with a Gaussian-smoothed target over bins (turning value regression into calibrated classification).
Setup.
Among our baselines, LSIQ (and SQIL) are natural candidates because their critics use MSE-like objectives, whereas IQ-Learn and RIZE optimize MaxEnt-style IRL objectives. We therefore apply the swap to LSIQ and keep all other components unchanged. Following LSIQ’s critic clipping in [200,200][-200,200], we set $v_{\min}=-200$, $v_{\max}=200$, use num_bins=101, and Gaussian width $\sigma=8$. We evaluate on Walker2d-v2, Ant-v2, and Hammer-v1 with three demonstrations.
Results.
As shown in Figure 14, HL-Gauss underperforms the original MSE-based LSIQ on all three tasks: it yields no improvement on Walker2d-v2 or Ant-v2, and remains near zero on Hammer-v1. A likely cause is structural: LSIQ’s critic objective is the sum of two squared-error terms (expert and policy), which is integral to its least-squares/mixture formulation; swapping those squared losses for a categorical (multi-bin) classification surrogate alters the optimization geometry and weakens the intended expert–policy mixture shaping. In contrast, standard Q-learning—where loss swaps have shown benefits—regresses to a single TD target, making the classification surrogate a closer drop-in for MSE.
Figure 14: Loss swap in LSIQ: MSE vs. HL-Gauss on Walker2d-v2, Ant-v2, and Hammer-v1 with three demonstrations. We keep $v_{\min}=-200$, $v_{\max}=200$, num_bins=101, and $\sigma=8$. HL-Gauss does not improve over MSE and remains near zero on Hammer-v1.
C.6 Ablations on Regularization Strategies
We study how the regularizer design affects performance on Walker2d-v2, Ant-v2, and Hammer-v1 with three demonstrations. Our baseline uses a squared TD-error regularizer $\Gamma$ in Equation (7) with separately optimized adaptive targets for expert and policy samples. We compare this to two alternatives: (i) a coupled target, where we replace the separate targets with a single shared $\lambda$ (initialized to 0 or $10$), and (ii) a plain L2-regularizer on rewards (no targets), akin to IQ-Learn’s constraint.
Figure 15 shows that squared TD with separate, adaptive targets yields the most robust and highest returns across tasks. Using a coupled target is brittle: $\lambda{=}10$ attains expert-like performance on Ant-v2, remains below the original result on Walker2d-v2, and fails on Hammer-v1, while $\lambda{=}0$ collapses across all tasks. Replacing squared TD with a plain L2 penalty also fails on all three tasks in our setup. We attribute this difference from IQ-Learn’s reports to implementation and hyperparameter choices tailored to our setting: RIZE employs an IQN critic, target policy networks, different learning rates (notably for the policy), a higher entropy coefficient (0.05 vs. 0.01), and a smaller regularizer coefficient (0.1 vs. 0.5). These choices were tuned for squared TD with adaptive targets; swapping to an L2 penalty disrupts learning. Overall, the results indicate that adaptive, decoupled targets are crucial for stabilizing reward across tasks.
Figure 15: Ablation of regularization strategies on selected tasks (Walker2d, Ant, Hammer) using three expert demonstrations. Results are expert-normalized and reported as the mean over five seeds with 95% confidence intervals.
C.7 Hyperparameter Tuning
We present our analysis and comparison of important hyperparameters utilized in our algorithm. Plots depict mean over five seeds with 95% confidence intervals.
Adaptive Targets.
@@ C.7 Hyperparameter Tuning
Selecting appropriate initial values and learning rates for the automatic fine-tuning of $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$ is critical in our approach. First, we observe that a suitable learning rate is essential for the stable training of our imitation learning agent, as illustrated in Figure 16(a). Our findings indicate that $\lambda^{\pi}$ must be optimized very slowly; using larger learning rates can destabilize training and hinder progress. In contrast, $\lambda^{\pi_{E}}$ demonstrates greater resilience when optimized with higher learning rates. Additionally, $\lambda^{\pi_{E}}$ remains robust even with varying initial values. However, as shown in Figure 16(b), failing to select an appropriate initial value for $\lambda^{\pi}$ can negatively impact learning. Overall, Figures 16(a) and 16(b) highlight the need for careful selection of both the learning rate and initial value when optimizing $\lambda^{\pi}$, while $\lambda^{\pi_{E}}$ exhibits considerable robustness in this regard.
@@
Figure 16: Fine-tuning analysis of adaptive targets. (a) Learning rates: Turquoise represents our method’s primary result with learning rates of $1\mathrm{e}{-4}$ for $\lambda^{\pi_{E}}$ and $1\mathrm{e}{-5}$ for $\lambda^{\pi}$. Orange and blue lines indicate higher learning rates (e.g., $1\mathrm{e}{-3}$) for $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$, respectively. (b) Starting values: Turquoise shows the main result with initial values of 10 for $\lambda^{\pi_{E}}$ and 5 for $\lambda^{\pi}$, while other lines explore different starting values. Three trajectories are used throughout the analysis.
Regularization Coefficient.
@@
Entropy Coefficient.
We observe that the entropy coefficient is a crucial hyperparameter in inverse reinforcement learning (IRL) problems. As shown in Figure 17(b), IRL methods typically require small values for $\alpha$, a point previously noted (Garg2021). With expert demonstrations available, an imitation learning (IL) policy does not need to explore for optimal actions, as these are provided by the demonstrations. Consequently, higher values of $\alpha$ can lead to training instability, ultimately resulting in policy collapse.